# Nonlinear Model Order Reduction of Dynamical Systems in Process Engineering: Review and Comparison

### v1

@@ Abstract
Computationally cheap yet accurate enough dynamical models are vital for real-time capable nonlinear optimization and model-based control. When given a computationally expensive high-order prediction model, a reduction to a lower-order simplified model can enable such real-time applications. Herein, we review state-of-the-art nonlinear model order reduction methods and provide a theoretical comparison of method properties. Additionally, we discuss both general-purpose methods and tailored approaches for (chemical) process systems and we identify similarities and differences between these methods. As manifold-Galerkin approaches currently do not account for inputs in the construction of the reduced state subspace, we extend these methods to dynamical systems with inputs. In a comparative case study, we apply eight established model order reduction methods to an air separation process model: POD-Galerkin, nonlinear-POD-Galerkin, manifold-Galerkin, dynamic mode decomposition, Koopman theory, manifold learning with latent predictor, compartment modeling, and model aggregation. Herein, we do not investigate hyperreduction (reduction of FLOPS). Based on our findings, we discuss strengths and weaknesses of the model order reduction methods.
Keywords: model simplification, hybrid model, invariant manifold, nonlinear Galerkin, machine learning
@@ 1.  Introduction
Process systems in chemical engineering are inherently nonlinear, e.g., as a result of reaction kinetics, transport phenomena, and thermodynamic relationships [1]. Over the past decades, detailed mechanistic process modeling has become increasingly common in (chemical) process systems engineering (PSE). In particular, mechanistic first-principles models, accounting for conservation of mass, energy, and (less often) momentum, along with constitutive relations, possess favorable generalization capabilities. Hence, “digital twins” and “digital shadows” have been established as a valuable tool for process monitoring and operator training [2, 3]. However, in such predictive models, the high number of nonlinear differential equations is usually computationally prohibitive for online optimization and model-based control methods such as nonlinear model predictive control (NMPC). Furthermore, process models are often stiff, making them even more challenging to handle [4].
@@
Model reduction is a popular technique to reduce the computational cost of optimization by means of model simplification. In this context, a reduction in the number of differential states is called “model order reduction” (MOR) and the resulting model is a “reduced-order model” (ROM) [5, 6]. Conversely, the original model is the “full-order model” (FOM). MOR is particularly used for spatially distributed systems, which are commonly described by partial differential equations or modeled as networks of lumped systems. Examples of high-order systems in PSE include computational fluid dynamics [7], distillation columns [8], heat exchangers [9], chemical reactors [10], and integrated process networks [11]. Fine-grained mechanistic modeling of such systems can result in thousands or millions of states.
@@
Although mechanistic models can be very high-dimensional, FOMs often exhibit a certain coherence characterized by a few dominant mechanisms or patterns, e.g., due to stiffness of the system [12]. Provided that a part of the transients vanish after a short period of time due to dissipative phenomena with a comparably fast exponential decay, we obtain a time-scale separation between fast stable decay and slow response [13]. As a result, the high-dimensional state evolves most of the time near a low-dimensional topological subspace, referred to as “slow manifold”, “inertial manifold”, or “invariant manifold” [14, 15]. An accurate ROM is obtained when finding a low-order approximation of the FOM on the slow manifold. Such reducible systems are also called “lumpable” [16].
@@
MOR methods construct low-order approximations of a FOM, i.e., a ROM, either by projecting the FOM onto the slow manifold [17, 18] or by identifying an empirical ROM directly from state data [19]. Projecting the FOM onto a linear subspace is called a “Galerkin projection” [20], a concept of high significance for this article. In many cases, a considerable reduction (removing $80\,\mathrm{\%}$ of the states or more) of a FOM is possible at a minor sacrifice of accuracy. However, the range of validity of a ROM is considerably reduced as well, e.g., highly transient situations, such as process startup or shutdown, are usually no longer covered.
@@
Reduction methods can be classified according to how the FOM is utilized [21]. Specifically, “intrusive methods” rely on structural and parametric knowledge of the FOM. A large subclass of intrusive methods are reduced-basis methods, which seek a low-dimensional set of basis vectors spanning a linear or affine subspace on which the system approximately evolves, followed by an (intrusive) projection of the FOM onto this subspace [22, 23]. Non-projection-based intrusive methods include simplifying assumptions and heuristic reduction. On the other hand, “non-intrusive methods” are fully detached from the FOM equations, i.e., model-free. These methods only require state time series data to generate a ROM.
@@
Many projection-based approaches follow a two-stage procedure to derive a ROM. First, the nonlinear system is reversibly transformed into a favorable representation, e.g., some canonical or normal form. Subsequently, reduction is performed on the transformed system by systematic truncation or residualization of dynamic components. Depending on the original model and the reduction method, the resulting ROM can be anything from a white-box (all equations/terms have a physical meaning) to an almost black-box (purely empirical) model. Examples of the first step include proper orthogonal decomposition [24], input-output balancing [25, 26], singular perturbation canonical form [27], Weierstrass canonical form [28], canonical bilinear or polynomial form [29, 30, 31], and Koopman lifting [32, 33].
@@
For some types of systems, e.g., convection-dominated systems, wave-type equations, and conservation problems, the assumption of a low-dimensional linear subspace does not hold [34]. Such systems live on a nonlinear submanifold rather than on a linear subspace [35]. Consequently, Galerkin projection requires a large number of basis vectors to retain high ROM accuracy [17, 36]. Instead, nonlinear subspace methods can approximate the nonlinear slow manifold more closely. Projecting the FOM onto the nonlinear submanifold is a “nonlinear Galerkin” [35] or “manifold-Galerkin” [36] projection. Fig. 1 illustrates the advantage of a nonlinear manifold. The state data generated by a 3D FOM approximately lie on a 2D linear (affine) or a 1D nonlinear subspace.
@@
When reducing a nonlinear FOM, the CPU cost of evaluating a ROM is often not significantly decreased, despite the lower differential order [22]. Thus, to achieve a reduction in computational costs, an additional consecutive “hyperreduction” step is performed [37, 38], which reduces the arithmetic complexity (FLOPS). Hyperreduction is achieved by applying supervised approximation techniques to replace expensive terms by, e.g., piecewise linear functions [39, 40], polynomials [41, 42, 43], sparse regression models [44], or artificial neural networks (ANNs) [45, 37]. In some cases, semi-empirical shortcut methods are an alternative to eliminate steady-state equations of a ROM [46]. Further, hyperreduction methods may also be implemented in an adaptive fashion [47, 48]. In case of a differential-algebraic ROM, hyperreduction automatically eliminates algebraic equations, so a prior reduction of algebraic variables [49, 50] is often not required. Combining data-driven and mechanistic elements, the resulting ROM can be considered a “hybrid model”. Hyperreduction drives the ROM further towards a black-box model but is essential for real-time capability. The overall model reduction process thus typically involves the two ingredients (i) MOR and (ii) hyperreduction.
@@
Although non-intrusive reduction methods do not operate on the FOM equations, these methods are still based on state information (and sometimes also on time derivative data). Conversely, the identification of a black-box model from input-output data does not involve any state data and thus lacks the information carried by the states. In particular, state information turn out to be highly valuable for many systems, where input-output system identification may struggle to extract sensible latent information [51].
@@
Despite a long history of research, MOR remains an evolving field with major contributions made in recent years, e.g., [19, 52, 33, 36]. Comprehensive reviews of MOR are found in [53, 18, 7, 54, 55] as well as textbooks on linear [56, 20] and nonlinear MOR [57, 58, 59]. However, despite an extensive body of literature, available low-barrier introductory texts [5, 6] and comparative studies in the PSE field [60, 61, 7] lack a discussion of some recent developments. Furthermore, reduced modeling strategies that have been developed in the chemical engineering community are often not discussed in the general model reduction literature.
@@
Over the years, managing the growing number of MOR options and making a suitable choice has become increasingly difficult. Therefore, this article aims to provide an overview of classical and more recent methods. We consider systems with inputs and focus on nonlinear state-space models, but we also include index-one differential-algebraic systems. We compare common methods both theoretically and in a PSE case study. As we target ROM application within NMPC and soft sensors, we want to preserve the knowledge of all states and retain flexibility in formulating the optimal control problem. Hence, we aim for ROMs that allow reconstruction of all original states and thus focus on input-to-state reduction methods rather than input-output methods. Consequently, we do not review input-output balancing methods for linear systems [25, 62], empirical extensions to nonlinear systems [63, 64, 26, 65], or data-driven moment matching [66]. Further, we disregard MOR methods based on linearized FOMs, e.g., applying feedback linearization before reduction [67].
@@
The manuscript is structured as follows. First, we introduce a general version of the MOR problem for systems with inputs in Section 2. In Section 3, we provide an overview of general-purpose MOR methods. We divide this review into intrusive methods in Section 3.1 and non-intrusive methods in Section 3.2. During the review of intrusive methods, we observe that current manifold learning methods only involve the system states and do not account for external inputs (Section 3.1.4). Hence, we propose an extension in Section 3.1.5. The review of general-purpose methods concludes with a theoretical comparison in Section 3.3. In Section 4, we review problem-specific MOR approaches that have been developed in the field of PSE and draw connections to methods in Section 3. Additionally, we review applications of MOR for three important unit operations, namely distillation columns, heat exchangers, and reactors. Section 5 provides a comparative case study, where we reduce the process model of an air separation unit (ASU). Therein, we compare eight MOR approaches that are selected as a representative ensemble of the methods reviewed. Notably, we limit ourselves to MOR and do not consider hyperreduction and CPU costs of integration to avoid an extra level of methods and complexity. However, we expect that any ROM of sufficiently low order can be made real-time capable by appropriate hyperreduction. In Section 6, we summarize and discuss the strengths and weaknesses of the MOR methods. Section 7 concludes the manuscript.
@@
Given a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$, $\mathbf{A}=[\bm{a}_{1},...,\bm{a}_{m}]$, the linear space spanned by the column vectors $\bm{a}_{i}\in\mathbb{R}^{n}$ is denoted by $\mathcal{A}=\mathrm{col}(\mathbf{A})$ and called the “column space” of $\mathbf{A}$. Given a real non-square matrix $\mathbf{A}$, the “pseudoinverse” is $\mathbf{A}^{+}\triangleq(\mathbf{A}^{\mathrm{T}}\mathbf{A})^{-1}\mathbf{A}^{\mathrm{T}}$. Given some domain $\mathcal{X}$, we denote by $\mathcal{C}^{r}(\mathcal{X})$ the set of $r$-times continuously differentiable functions over $\mathcal{X}$, $r\geq 1$, and call a function $f$ smooth if $f\in\mathcal{C}^{\infty}(\mathcal{D})$. We denote the identity matrix by $\mathbf{I}$. Given a single independent coordinate $t\in\mathbb{R}$ (here time), we call the map $\bm{\gamma}:\mathbb{R}\rightarrow\mathbb{R}^{n}$ a “curve”. We denote the velocity of a curve $\bm{\gamma}$ at time $t$ by $\dot{\bm{\gamma}}(t)\triangleq\frac{\mathrm{d}\gamma}{\mathrm{d}t}|_{t}$. Given some $\mathcal{C}^{r}$-map $\bm{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ between Euclidean spaces $\mathbb{R}^{n}$ and $\mathbb{R}^{m}$, where $n,m\geq 1$, we denote the linear “differential map” of $\bm{f}$ at $\bm{x}\in\mathbb{R}^{n}$ by $\mathrm{D}\bm{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$. Given some point $\bm{x}\in\mathbb{R}^{n}$, we associate $\mathrm{D}\bm{f}$ with its Jacobian matrix $\mathrm{D}\bm{f}(\bm{x})\in\mathbb{R}^{m\times n}$. For a function $\bm{f}(\bm{x},\bm{y})$ of two arguments, we consider $\mathrm{D}_{\bm{x}}\bm{f}(\bm{x},\bm{y})$ and $\mathrm{D}_{\bm{y}}\bm{f}(\bm{x},\bm{y})$ separately. An “immersion” is a smooth map, $\bm{h}:\mathcal{Z}\subseteq\mathbb{R}^{n_{z}}\rightarrow\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$, satisfying $\mathrm{rank}(\mathrm{D}_{z}\bm{h}^{\dagger}(\,\cdot\,))=n_{z}$ (injective) everywhere in $\mathcal{Z}$.
@@
=$\mathcal{M}=\{\bm{x}\in\mathbb{R}^{n}:\bm{c}(\bm{x})=\bm{0}\}\,.$ (1)
Assume that $\mathcal{M}$ is non-empty and the Jacobian $\mathrm{D}\bm{c}(\bm{x})\in\mathbb{R}^{m\times n}$ has constant rank $m$ at all $\bm{x}\in\mathcal{M}$. Then, $\mathcal{M}$ is a manifold of class $\mathcal{C}^{r}$, i.e., a differentiable manifold. Moreover, $\mathcal{M}$ has dimension $k=n-m$ and is embedded in $\mathbb{R}^{n}$, and therefore called an “embedded $k$-submanifold” of $\mathbb{R}^{n}$. By the definition of continuity, the preimage $\mathcal{M}=\bm{c}^{-1}(\{\bm{0}\})$ is closed in $\mathbb{R}^{n}$ and therefore $\mathcal{M}$ is said to be properly embedded.
We remark that the embedding is not an essential part of the definition, i.e., there exist more generic definitions that are independent of an ambient space [68]. Therefore, we sometimes drop the terms “embedded” or “sub” (manifold). Note that $\mathbb{R}^{n}$ is a manifold itself. By limiting ourselves to embedded submanifolds of (open subsets $\mathcal{X}\subseteq\mathbb{R}^{n}$ of) Euclidean space $\mathbb{R}^{n}$ defined through level sets, we circumvent geometric concepts like charts and tangent bundles. We refer the interested reader to [68, 69, 70].

### v2

@@ Abstract
Computationally cheap yet accurate dynamical models are a key requirement for real-time capable nonlinear optimization and model-based control. When given a computationally expensive high-order prediction model, a reduction to a lower-order simplified model can enable such real-time applications. Herein, we review nonlinear model order reduction methods and provide a comparison of method characteristics. Additionally, we discuss both general-purpose methods and tailored approaches for chemical process systems and we identify similarities and differences between these methods. As machine learning manifold-Galerkin approaches currently do not account for inputs in the construction of the reduced state subspace, we extend these methods to dynamical systems with inputs. In a comparative case study, we apply eight established model order reduction methods to an air separation process model: POD-Galerkin, nonlinear-POD-Galerkin, manifold-Galerkin, dynamic mode decomposition, Koopman theory, manifold learning with latent predictor, compartment modeling, and model aggregation. Herein, we do not investigate hyperreduction, i.e., reduction of floating point operations. Based on our findings, we discuss strengths and weaknesses of the model order reduction methods.
Keywords: model simplification, hybrid model, invariant manifold, nonlinear Galerkin, non-autonomous system
@@ 1. Introduction
Process systems in chemical engineering are inherently nonlinear, e.g., as a result of reaction kinetics, transport phenomena, and thermodynamic relationships [176]. Over the past decades, detailed mechanistic process modeling has become increasingly common in (chemical) process systems engineering. In particular, mechanistic first-principles models, accounting for conservation of mass, energy, and (less often) momentum, along with constitutive relations, possess favorable generalization capabilities. Hence, digital twins and digital shadows have been established as a valuable tool for process monitoring and operator training [213, 139]. However, in such predictive models, the high number of nonlinear differential equations is usually computationally prohibitive for online optimization and model-based control methods such as nonlinear model predictive control. Furthermore, process models are often stiff, making them even more challenging to handle [33].
Table 1: List of acronyms used in this work.
Acronym Term
ANN Artificial neural network
AEN Autoencoder network
AGG Aggregation method
COMP Compartment method
DAE Differential-algebraic system of equations
DMD Dynamic mode decomposition
DMDc DMD with controls (or general inputs)
FOM Full-order model
HX Heat exchanger
KW Koopman-Wiener model
MFL Manifold learning
MFL-ANN MFL combined with latent ANN predictor
MFLu Manifold learning with inputs
MOR Model order reduction
ODE Ordinary differential system of equations
POD Proper orthogonal decomposition
POD-Res POD residualization
ROM Reduced order model
RMSE Root mean squared error
SPT Singular perturbation theory
SVD Singular value decomposition
@@
Model reduction is a popular technique to reduce the computational cost of optimization by means of model simplification. In this context, a reduction in the number of differential states is called model order reduction (MOR) and the resulting model is a reduced-order model (ROM) [186, 4]. Conversely, the original model is the full-order model (FOM). MOR is particularly used for spatially distributed systems, which are commonly described by partial differential equations or modeled as networks of lumped systems. Examples of high-order process systems include computational fluid dynamics [249], distillation columns [299], heat exchangers [101], chemical reactors [80], and integrated process networks [55]. Fine-grained mechanistic modeling of such systems can result in thousands or millions of states.
@@
Although mechanistic models can be very high-dimensional, FOMs often exhibit a certain coherence characterized by a few dominant mechanisms or patterns, e.g., due to stiffness of the system [207]. Provided that a part of the transients vanish after a short period of time due to dissipative phenomena with a comparably fast exponential decay, we obtain a time-scale separation between fast stable decay and slow response [92]. As a result, the high-dimensional state evolves most of the time near a low-dimensional topological subspace, referred to as slow manifold [203]. Because this subspace is characterized by a set of invariance equations, i.e., the (quasi-)steady-state conditions of the fast stable phenomena, the slow manifold is a special type of invariant manifold [68, 241, 151]. An accurate ROM is obtained when finding a low-order approximation of the FOM on the slow manifold. Such reducible systems are also called lumpable [124].
@@
MOR methods construct low-order approximations of a FOM, i.e., a ROM, either by projecting the FOM onto the slow manifold [269, 18] or by identifying an empirical ROM directly from state data [257, 148]. Projecting the FOM onto a linear subspace is called a Galerkin projection [5], a concept of high significance for this article. Below, we also discuss nonlinear variants of Galerkin methods as well as other projection methods. In many cases, a considerable reduction (removing $80\,\mathrm{\%}$ of the states or more) of a FOM is possible at a minor sacrifice of accuracy. However, the range of validity of a ROM is considerably reduced as well, e.g., highly transient situations, such as process startup or shutdown, are often no longer covered.
@@
Reduction methods can be classified according to how the FOM is utilized [251]. Specifically, intrusive methods rely on structural and parametric knowledge of the FOM. Note that the term “intrusive” as adopted here is fully independent from its use in structural mechanics [193], and as such partially inconsistent. A large subclass of intrusive methods are reduced-basis methods, which seek a low-dimensional set of basis vectors spanning a linear or affine subspace on which the system approximately evolves, followed by an (intrusive) projection of the FOM onto this subspace [235, 206]. Non-projection-based intrusive methods include simplifying assumptions and heuristic reduction. On the other hand, non-intrusive methods are fully detached from the FOM equations, i.e., model-free. These methods only require time series data of the model states to generate a ROM.
@@
Many projection-based approaches follow a two-stage procedure to derive a ROM. First, the nonlinear system is reversibly transformed into a favorable representation, e.g., some canonical or normal form [195, 267, 242]. Subsequently, reduction is performed on the transformed system by systematic truncation or residualization of dynamic components. Depending on the original model and the reduction method, the resulting ROM can be anything from a white-box (all equations have a physical meaning) to an almost black-box (purely empirical) model. Examples of the first step include proper orthogonal decomposition [25], nonlinear normal modes [267, 288], input-output balancing [195, 98], singular perturbation canonical form [140], Weierstrass canonical form [245], canonical bilinear or polynomial form [226, 66, 95], and Koopman lifting [191, 298].
@@
For some types of systems, e.g., convection-dominated systems and wave-type equations, the assumption of a low-dimensional linear subspace does not hold [15, 117]. Such systems live on a nonlinear submanifold rather than on a linear subspace [131, 267]. Consequently, Galerkin projection requires a large number of basis vectors to retain high ROM accuracy [269, 159]. Instead, nonlinear subspace methods can approximate the nonlinear slow manifold more closely. Projecting the FOM onto the nonlinear submanifold is a nonlinear Galerkin [131] or manifold-Galerkin [159] projection, which is closely connected to the concept of nonlinear normal modes in mechanics [267]. Figure 1 illustrates the advantage of using a nonlinear manifold as the reduced subspace. Although the dimensionality of the 1D nonlinear subspace is lower than the 2D linear subspace, the data points generated by the 3D full-order system are well captured by either submanifold. Consequently, using a nonlinear subspace method may enable a similarly accurate ROM of lower order.
@@
When reducing a nonlinear FOM, the CPU cost of evaluating a ROM is often not significantly decreased, despite the lower differential order [235]. Thus, to achieve a reduction in computational costs, an additional consecutive hyperreduction step is performed [100, 252], which reduces the arithmetic complexity, i.e., floating point operations per model call. Hyperreduction is achieved by applying supervised approximation techniques to replace expensive terms by, e.g., piecewise linear functions [17, 238], polynomials [30, 201, 47], sparse regression models [35], or artificial neural networks (ANNs) [270, 100]. In some cases, semi-empirical shortcut methods are an alternative to eliminate steady-state equations of a ROM [58]. Further, hyperreduction methods may also be implemented in an adaptive fashion [114, 175]. In case of a differential-algebraic ROM, hyperreduction automatically eliminates algebraic equations, so a prior reduction of algebraic variables [160, 275] is often not required. Combining data-driven and mechanistic elements, the resulting ROM can be considered a hybrid model. Hyperreduction drives the ROM further towards a black-box model but is essential for real-time capability. The overall model reduction process thus typically involves the two ingredients (i) MOR and (ii) hyperreduction.
@@
Although non-intrusive reduction methods do not operate on the FOM equations, these methods are still based on FOM state information (and sometimes also on time derivative data). Conversely, the identification of a low-order black-box model from input-output data (e.g., using subspace identification [232], Loewener framework [189, 218], or AAA algorithm [198, 243, 32]) does not involve the original FOM states and thus lacks the full information carried by the FOM states. In particular, FOM state information turns out to be highly valuable in some MOR problems, where input-output system identification may fail to extract sensible latent information [261].
@@
Despite a long history of research, MOR remains an evolving field with major contributions made in recent years, e.g., [257, 9, 298, 159]. Comprehensive reviews of MOR are found in [84, 18, 249, 155, 183] as well as textbooks on linear [205, 5] and nonlinear MOR [204, 122, 23, 280]. However, although an extensive body of literature is available, low-barrier introductory texts [186, 4] and comparative studies in the process systems field [256, 99, 249] lack a discussion of some recent developments. Furthermore, reduced modeling strategies that have been developed in the chemical engineering community are often not discussed in the general model reduction literature.
@@
Over the years, managing the growing number of MOR options and making a suitable choice has become increasingly difficult. Therefore, this article aims to provide an overview of classical and more recent methods. We consider systems with inputs and focus on nonlinear state-space models, but we also include index-one differential-algebraic systems. We compare common methods both theoretically and in an industrially relevant case study.
As we target ROM application within numerical optimization and model predictive control, we want to preserve the knowledge of all states and retain flexibility in formulating the optimal control problem. Hence, we aim for ROMs that allow reconstruction of all original states and thus focus on input-to-state reduction methods rather than input-output methods. Consequently, we do not review input-output balancing methods for linear systems [195, 223], empirical balancing of nonlinear systems [255, 156, 98, 119], or data-driven moment matching [253]. However, we remark that for problems with fixed input-output configuration, balancing methods can provide very low-order ROMs of high accuracy and guaranteed stability preservation [138]. Further, we disregard MOR methods based on linearized FOMs, e.g., applying feedback linearization before reduction [11]. Finally, we remark that model reduction of non-autonomous systems with inputs shares some conceptual similarities with MOR of stochastic dynamical systems [242, 59]. Specifically, both systems feature a-priori unknown external variables. However, we do not investigate these connections further.
@@
The manuscript is structured as follows. First, we present a general version of the MOR problem for systems with inputs in Section 2. Therein, we characterize the reduced subspace based on the invariance equations and explicitly account for system inputs. In Section 3, we provide an overview of general-purpose MOR methods. We divide this review into intrusive methods in Section 3.1 and non-intrusive methods in Section 3.2. During the review of intrusive methods, in Section 3.1.4, we observe that current manifold learning methods do not account for external inputs. Hence, we propose an extension of the machine learning manifold-Galerkin method to non-autonomous systems with inputs in Section 3.1.5. The review of general-purpose methods concludes with a theoretical comparison in Section 3.3. In Section 4, we review problem-specific MOR approaches that have been developed in the field of process systems engineering and draw connections to methods in Section 3. Additionally, we review applications of MOR for three important unit operations, namely distillation columns, heat exchangers, and reactors. Section 5 provides a comparative case study, where we reduce the process model of an air separation unit. Therein, we compare eight MOR approaches that are selected as a representative ensemble of the methods reviewed. Notably, we limit ourselves to MOR and do not consider hyperreduction and CPU costs of integration to avoid an extra level of methods and complexity. However, we expect that any ROM of sufficiently low order can be made real-time capable by appropriate hyperreduction. In Section 6, we summarize and discuss the strengths and weaknesses of the MOR methods. Section 7 concludes the manuscript.
@@
Given a matrix $\mathbf{A}\in\mathbb{R}^{n\times m}$, $\mathbf{A}=[\bm{a}_{1},...,\bm{a}_{m}]$, the linear space spanned by the column vectors $\bm{a}_{i}\in\mathbb{R}^{n}$ is denoted by $\mathcal{A}=\mathrm{col}(\mathbf{A})$ and called the column space of $\mathbf{A}$. For brevity, we prefer this notation over $\mathcal{A}=\mathrm{span}\{\bm{a}_{1},...,\bm{a}_{m}\}$. Given a real non-square matrix $\mathbf{A}$, the pseudoinverse is $\mathbf{A}^{+}\triangleq(\mathbf{A}^{\mathrm{T}}\mathbf{A})^{-1}\mathbf{A}^{\mathrm{T}}$. Given some domain $\mathcal{X}$, we denote by $\mathcal{C}^{r}(\mathcal{X})$ the set of $r$-times continuously differentiable functions over $\mathcal{X}$, $r\geq 1$, and call a function $f$ smooth if $f\in\mathcal{C}^{\infty}(\mathcal{D})$. We denote the identity matrix by $\mathbf{I}$ and the identity map by $\mathbf{id}$. Given a single independent coordinate $t\in\mathbb{R}$ (here time), we call the map $\bm{\gamma}:\mathbb{R}\rightarrow\mathbb{R}^{n}$ a curve. We denote the velocity of a curve $\bm{\gamma}$ at time $t$ by $\dot{\bm{\gamma}}(t)\triangleq\frac{\mathrm{d}\gamma}{\mathrm{d}t}|_{t}$. Given some $\mathcal{C}^{r}$-map $\bm{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ between Euclidean spaces $\mathbb{R}^{n}$ and $\mathbb{R}^{m}$, where $n,m\geq 1$, we denote the linear differential map of $\bm{f}$ at $\bm{x}\in\mathbb{R}^{n}$ by $\mathrm{D}\bm{f}{(\bm{x})}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$. Given some point $\bm{x}\in\mathbb{R}^{n}$, we associate $\mathrm{D}\bm{f}{(\bm{x})}$ with its Jacobian matrix $\mathrm{D}\bm{f}(\bm{x})\in\mathbb{R}^{m\times n}$. For a function $\bm{f}(\bm{x},\bm{y})$ of two arguments, we consider $\mathrm{D}_{\bm{x}}\bm{f}(\bm{x},\bm{y})$ and $\mathrm{D}_{\bm{y}}\bm{f}(\bm{x},\bm{y})$ separately. An immersion is a smooth map, $\bm{h}:\mathcal{Z}\subseteq\mathbb{R}^{n_{z}}\rightarrow\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$, satisfying $\mathrm{rank}(\mathrm{D}_{z}\bm{h}^{\dagger}(\,\cdot\,))=n_{z}$ (injective) everywhere in $\mathcal{Z}$.
@@
$\mathcal{M}=\{\bm{x}\in\mathbb{R}^{n}:\bm{c}(\bm{x})=\bm{0}\}\,.$ (1)
Assume that $\mathcal{M}$ is non-empty and the Jacobian $\mathrm{D}\bm{c}(\bm{x})\in\mathbb{R}^{m\times n}$ has constant rank $m$ at all $\bm{x}\in\mathcal{M}$. Then, $\mathcal{M}$ is a manifold of class $\mathcal{C}^{r}$, i.e., a differentiable manifold. Moreover, $\mathcal{M}$ has dimension k=nmk=n-m and is embedded in $\mathbb{R}^{n}$, and therefore called an embedded $k$-submanifold of $\mathbb{R}^{n}$. By the definition of continuity, the preimage $\mathcal{M}=\bm{c}^{-1}(\{\bm{0}\})$ is closed in $\mathbb{R}^{n}$ and therefore $\mathcal{M}$ is said to be properly embedded.
We remark that the embedding is not an essential part of the definition, i.e., there exist more generic definitions that are independent of an ambient space [158]. Therefore, we sometimes drop the terms embedded or “sub” (manifold). Note that $\mathbb{R}^{n}$ is a manifold itself. By limiting ourselves to embedded submanifolds of open subsets $\mathcal{X}\subseteq\mathbb{R}^{n}$ of Euclidean space $\mathbb{R}^{n}$ defined through level sets, we circumvent geometric concepts like charts and tangent bundles. We refer the interested reader to [158, 197, 31, 37]. Below, we always consider the case $m<n$, where $n=n_{x}$ (FOM dimension) and $k=n_{z}\geq 1$ (ROM dimension). For notational and conceptual simplicity, we will focus on $\mathcal{C}^{\infty}$ maps and $\mathcal{C}^{\infty}$ manifolds. However, we remark that most methods are applicable to a less restrictive setup, such as $\mathcal{C}^{2}$, which is occasionally also referred to as “smooth” in the literature.